What is the value of ni for an electron that emits a photon of wavelength 93.85 nm when it returns to the ground state in the H atom?
1
Expert's answer
2020-02-18T06:38:56-0500

It is given that electrons emits a photon of wavelength 93.85 nm and then returns to ground state from ni the energy state, "\lambda = 93.85 \times 10^-7 cm"

now using Rydberg formula for wavelength of photon ejected

we know that

"\frac{1}{\lambda}" ="R_h(\frac{1}{n_f^2} -\frac{1}{n_i^2})"

here, "\lambda" = wavelength of photon ejected, "n_i =" energy level from which electron coming and "n_f" = energy level at which it is going=1

"R_h" = Rydberg constant = 109737.32 cm-1

now ,

"\frac{1}{\lambda} = 109737.32 ( \frac{1}{1^2} - \frac{1}{n_i^2} )"


"\frac{1}{93.85 \times10^-7} = 109737.32 (1- \frac{1}{n_i^2})"


.010655 "\times 10^7 =109737.32 (1-\frac{1}{n_i^2})"


"\frac{1.06 \times10^5}{109737.32 } =(1- \frac{1}{n_i^2})"


.97098= 1 "-\frac{1}{n_i^2}"

"\frac{1}{n_i^2} = 1-.97098"


"n_i^2 = \frac{1}{.029}"


"n_i^2= 34.46"


"n_i" =5.87

so here we take minimum energy level as "n_i" =6